# Investigation about a statement equivalent to Riemann Hypothesis (RH) applied to Dirichlet primitive L functions

### v15

@@ 1. Introduction
Using Dirichlet characters $\chi(n)$, L functions, [3, p. 249, and 262], and, specially, ([6, p. 3, and 37] (whose conventions, i.e. $\chi_{0}$ as principal, is used throughout) are defined as :
@@
For reasons will be clear below we focus only on $\chi_{primitive}$, defined in [3, p. 168 -170]. Companion function like $\xi(s)$, ( [6, p. 62] ) for $\zeta(s)$, is $\xi(s,\chi_{primitive})$, ( [6, p. 71] ) for $L(s,\chi_{primitive})$, below their definition :
@@
where $\Gamma(z)$ is defined in [3, p. 251], $q=congruence\quad modulus$. While $\alpha=0$ if $\chi_{primitive}(-1)=1$, even character, and, $\alpha=1$ if $\chi_{primitive}(-1)=-1$, odd character. Both (1.2 ) have the same zeros of the $\zeta(s)$ or $L(s,\chi)$ functions respectively in the critical strip [6] [7, p. 16]. Following [6, p. 71, and 79] the functional equation for $\xi(s)$ and $\xi(s,\chi)$ are:
@@
Where $\bar{\chi}$ is the complex conjugate of $\chi$ and $\tau(\chi)$ is the gauss sum $\tau(\chi):=\sum_{m=1}^{q}\chi(m)e^{2\pi im/q}$ see [6, p. 65] , or [3, p. 165].
@@
  • In appendix B we compare an useful approximation to (3.13 ).
@@ 2.1. LEMMA 1
Let us apply Lemma 1 to 1.3 ( remember $\chi_{primitive}$). For [6, p. 66] $|\tau(\chi)|^{2}=q$, then $\left|\frac{i^{\alpha}q^{\frac{1}{2}}}{\tau(\chi)}\right|=1$. But, at $\epsilon=0$ we have: $1-s=\bar{s}$. Besides, as $\Gamma(\bar{s})=\int_{0}^{\infty}x^{\bar{s}-1}e^{-x}dx=\bar{\Gamma}(s)$ ([7, p. 8]), $\left(\frac{q}{\pi}\right)^{\frac{\bar{s}+\alpha}{2}}=\bar{\left(\frac{q}{\pi}\right)^{\frac{s+\alpha}{2}}}$, and $L(\bar{s},\bar{\chi})=\bar{L}(s,\chi)$ ( (1.1)) , then for 1.2 ), and ( 1.3):
@@ 3. Lemmas on $L(s,\chi_{primitive})$ phase variations along $t$ computed by Euler product
If $\chi$ is not principal, as a primitive character, then $L(s,\chi)$ is an entire function ([3, p. 255] ), and, (1.1) converges absolutely for $\Re(s)>1$, while converges conditionally (see [1, p. 406] ) for $\Re(s)>0$ ( [10, p. 7], and, also unifomily [5, p. 7]). So we take $\Re[s]>0$ throughout.
@@ 3.1. Euler product in critical strip
For example for $p_{max}=31$ we have in (3.2 ) that $N^{o}(terms)=787320$, whoose only 31 belong to $L(s,\chi,p_{max})$. The others belong to $R(s,\chi,p_{max})=R(t,\epsilon,\chi,p_{max})$.
So we does not need to care, in applying Euler rearrangement for $L(s,\chi,p_{max})=\sum_{n=1}^{p_{max}}\frac{\chi(n)}{n^{s}}$, about absolute convergence of the $\prod_{\forall p}\frac{1}{1-\frac{\chi(p)}{p^{s}}}\quad p\quad prime$, because we have an huge, but, finite amount of terms.
@@
Only first term in right part of (3.1), i.e $L(s,\chi,p_{max})=\sum_{n=1}^{p_{max}}\frac{\chi(n)}{n^{s}}$, tend to (C.1), so, due to (3.3), we have, considering the whole critical strip:
@@
But if we consider the difference (or better the phase difference) in a finite interval, suitably far from $\L (s.\chi)=0$, the conclusion is different. Let us take:
@@
where in (3.3) we have an huge but finite number of terms with zero mean value. If $p_{max}>>p^{*}$ we have a lot of cancellation in $\Delta t$ interval for single (3.6) addend. The cancellation will be complete, for a particular $\rho$, only if
(3.7) $\Delta t\times\ln(\rho)=\frac{2\pi}{\ln(p^{*})}\ln(\rho)=m\times 2\pi\ ;\ m\in N\ ,\ i.e.\ \rho=(p^{*})^{m}$
@@
Such $\Delta t_{\rho}$ become smaller in absolute value, and, also the amplitude of oscillation evidenced in (3.3), decreases as $p_{max}$ ( and, cosequently $\rho$ ) grows.
So the error terms at $\Delta t$ extrema must go to zero as $p_{max}\rightarrow\infty$, even if extremely slowly, because all $\ln(\rho)\times t$ are different.
@@
Let us consider $(R(t,\epsilon,\chi,p_{max})$ terms with $\rho=p_{1}^{\alpha_{1}}...p_{max}^{\alpha_{max}}=p_{max}\times r$, and $\frac{\ln(r)}{\ln(p_{max})}$ small. These terms give an almost null contribution to the differences from $t_{2}$ to $t_{1}$ of (3.1 ), i.e. $L(s,\chi,p_{max})+R(s,\chi,p_{max})$ for $n$ and $\rho$ close to $p_{max}$ because phase difference in $\Delta t$ is: $\ln(\rho)\Delta t=\ln(p_{max})\left(1+\frac{\ln(r)}{\ln(p_{max})}\right)\ \frac{2\pi}{\ln(p_{max})}=2\pi\left(1+\frac{\ln(r)}{\ln(p_{max})}\right)$ , and for $p_{max}\rightarrow\infty$ it tends to $2\pi$. So as concern these small $r$ terms we have (considering mean values in $\Delta t$ interval):
@@
where $p^{*}$ is whatever big.
So condition (3.8) is an hint to convergence at lower $p_{max}$ while the argument stemming from identity (3.6) can be applied to generic $p_{max}>>p^{*}$ condition, but, it requires higher $p_{max}$ values to reach a comparable convergence. In a simplistic view we could say:
(3.3) is mainly an oscillation with period close to $\frac{2\pi}{\ln(p_{max})}$, if we take $\Delta t=m\times\frac{2\pi}{\ln(p_{max})}\ ;\ m\in N$ we filter away the main part of (3.3) .
@@ 3.2. Application of Euler Product in critical strip
Expression $\angle[L(s.\chi_{primitive})]=\angle\left[\sum_{n=1}^{\infty}\ \chi(n)/n^{-1}\right]$, where total phase is involved, converges for $\Re(s)>0$, but, in singular points.They are, see (2 ) , $s^{*}_{k}=1/2+\epsilon^{*}_{k}+it_{k}^{*}$,where $\xi(s^{*}_{k},\chi_{primitive})=L(s^{*}_{k},\chi_{primitive})=0$ (see (1.2 ), and (A.2) ). We neglect, for the moment, double zeros. To be specific $\xi(s,\chi)$ is equal ( [6, p. 82]) to the infinite Hadamard product, that has single factors like
@@
Now let us see how to evaluate phase variations of expression $\angle[L(s.\chi_{primitive})]=\angle\left[\sum_{n=1}^{\infty}\ \chi(n)/n^{-1}\right]$ using (3.9 ).
@@
It happens that , for $\{p_{max}\rightarrow\infty\}p^{*}\rightarrow\infty]$, (3.13) grows without bounds not only on the zeros, but also on the segment $0.5<\epsilon<\epsilon^{*}_{k}\ ,\ t=t_{k}^{*}$, where a step discontinuity arises in total phase of (3.11). So, when computing (3.13), we find a kind of ridge stemming from the $k^{th}$ zero $s^{*}_{k}=1/2+\epsilon^{*}_{k}+it_{k}^{*}\ \ ,\ 0.5>\epsilon^{*}_{k}\geq 0$ toward negative reals. Notice that for $p$ big ( i.e. $p^{1/2+\epsilon}>>1\geq\left|\cos(\ln(p)t-\angle[\chi(p)])\right|\ \ \$), the content of braces summed in (3.13 tends to be:
@@
We want to point out at $h$-prime classes in (3.13). If we limit to a finite $p_{max}$ we can write (3.13 ) as the sum of the $\phi(q)$ sums involving the $h$-prime classes:
@@ 3.4. LEMMA 5: Euler Product on critical line
PROOF. From Lemma 2, appendix A and (3.13 ) follows (3.22 ). End of proof.
@@
  • (3) $\frac{\Delta\angle[L_{EP}(t,\epsilon,\chi)]}{\Delta t}$ instead is defined everywhere but points $(\epsilon\leq\epsilon^{*}_{k}\ ,t=t^{*}_{k})$ , see (3.13 ). Here is used total phase.
@@
Suppose $\epsilon>0$ but arbitrarily small, $|t-t^{*}_{k}|>\frac{2\pi}{\ln(p^{*})}$, with (3.5), and $p^{*}$ big (fixed), and $p_{max}\rightarrow\infty$. Then,for (3.22) we have that (3.13) evaluated in $\epsilon=0$ is tending to $\ln\left[\sqrt{\frac{tq}{2\pi}}\right]<0$ as $(p_{max}\rightarrow\infty)p^{*}\rightarrow\infty]$’, in order to verify Lemma 2. We affirm that the same (3.13), evaluated for small $\epsilon>0$ is always negative but with lower absolute value, so that (3.22 ) has a positive increment for small $\epsilon>0$. i,e (at least for odd primitive characters, see (3.21) ):
@@
Negative contribution of (3.13) (and of final result) is given by the $h$-class primes (weighted by (3.26 ) function) inside the analogous intervals $\Delta y_{k,h}$ ( that for $k$ big tends to the previous one $\Delta x_{k,h}$):
@@
Positive contribution of (3.13) (and of final result) is given by the $h$-class primes (weighted by (3.26 ) function) inside the analogous intervals $\Delta y_{k,h}^{\prime}$ ( that for $k$ big tends to the previous one $\Delta x_{k,h}^{\prime}$)
@@
if $m$ is even, the factor $\sin\left(\frac{\pi\ln(x)}{\ln(p^{*})}\right)$ in (3.19) is positive, so, the contributions in corresponding $x$-intervals are as above. In other words, the sign contributions of final result, i.e. (3.13 ) minus (3.19), or , (3.20), in intervals $\Delta x,\Delta x^{\prime},\Delta y,\Delta y^{\prime}$, is as that stated in ( 3.29), (3.32) , and, (3.28), (3.31). Instead where $m$, in (3.33), is odd, then, in every interval above the sign changes concurrently. There are also intermediate situations, but, we will see they can be neglected. So we could compute (3.13) minus (3.19), till $p_{max}$, ( interchanging $O^{+}\leftrightarrow O^{-}$ when $m$ is odd in (3.33) ) as in :
@@
We stress again, (3.34 ) is computed without changing the order given by $p$ or $x=y$, like in (3.18 ) , i.e. partial sums are unaffected. Limit difference in (3.34 ) is in form: $\infty$-$\infty$, and, we know that if $|t-t_{k}^{*}|>\frac{2\pi}{\ln(p^{*})}$ limit result is finite . In other words: positive quantities: $O^{+}_{kh}(\epsilon)_{\int dLi}\ \ ,\ \ O^{-}_{k,h}(\epsilon)_{\int dLi}$,and , $O^{+}_{k,h}(\epsilon)_{\sum}\ \ ,\ \ O^{-}_{k,h}(\epsilon)_{\sum}$ like in ( 3.29), and (3.32) , and, (3.28), (3.31) are only “place-holders” filled according to progress of $x=y<p_{max}$, and with chosen sign following the $m$ parity in (3.33). Consider now the ratios :
@@
for each $h$ and for $k$ big, defined respectively in intervals $\Delta y_{k,h}\rightarrow\Delta x_{k,h}$ ( see (3.28) and (3.29 ) ) , and in intervals $\Delta y_{k,h}^{\prime}\rightarrow\Delta x_{k,h}^{\prime}$ ( see (3.31) and (3.32) ) .
LEMMA 7: $\forall h,k$ as $k\rightarrow\infty$ both (3.37 ) tend to unity.
PROOF. PNT ratio in small intervals [8, p. 23] assures that:
@@
i.e. for $x$ big enough the distribution of primes in equal-sign intevals $\Delta x_{k,h}\rightarrow\Delta y_{k,h}$, and, for (3.16) also $h$-class primes, can be close to theorethical one as we want. In [17, p. 65] it is dealt about the sum of prime sampled functions: “in the neighborhood of the number $x$ the average density of the primes is $1/\ln(x)$. On this basis, if one should wish an estimate for the sum of $F(p)$ over all primes $p<x$, the natural approximation would be”
@@
Now we want the ratio of all positive contributions to final result (3.34 ) at different $\epsilon$ value. In detail : $\forall h<q\ ,gcd(h,q)=1$, $\epsilon^{\prime\prime}>0$, and $\epsilon^{\prime}=0$:
@@
(3.44 ) and (3.43 ), if $\epsilon>0$ is fixed and $p_{max}\rightarrow\infty$ goes to zero, because if $\epsilon>0$ and $k$ i.e. $p_{max}\rightarrow\infty$, then $p^{-\epsilon}\rightarrow 0$. But we have $\forall p_{max}$:
@@
Besides the indeterminate limit (3.34 ), of $\infty$-$\infty$ form, for $\epsilon>0$ ($p^{*}$ is big and fixed), gives :

### v16

@@ 1. Introduction
Using Dirichlet characters $\chi(n)$, L functions, [3, p. 249, and 262], and, specially, ([7, p. 3, and 37] (whose conventions, i.e. $\chi_{0}$ as principal, is used throughout) are defined as :
@@
For reasons will be clear below we focus only on $\chi_{primitive}$, defined in [3, p. 168 -170]. Companion function like $\xi(s)$, ( [7, p. 62] ) for $\zeta(s)$, is $\xi(s,\chi_{primitive})$, ( [7, p. 71] ) for $L(s,\chi_{primitive})$, below their definition :
@@
where $\Gamma(z)$ is defined in [3, p. 251], $q=congruence\quad modulus$. While $\alpha=0$ if $\chi_{primitive}(-1)=1$, even character, and, $\alpha=1$ if $\chi_{primitive}(-1)=-1$, odd character. Both (1.2 ) have the same zeros of the $\zeta(s)$ or $L(s,\chi)$ functions respectively in the critical strip [7] [8, p. 16]. Following [7, p. 71, and 79] the functional equation for $\xi(s)$ and $\xi(s,\chi)$ are:
@@
Where $\bar{\chi}$ is the complex conjugate of $\chi$ and $\tau(\chi)$ is the gauss sum $\tau(\chi):=\sum_{m=1}^{q}\chi(m)e^{2\pi im/q}$ see [7, p. 65] , or [3, p. 165].
@@
  • In appendix B we compare an useful approximation to (3.16 ).
@@ 2.1. LEMMA 1
Let us apply Lemma 1 to 1.3 ( remember $\chi_{primitive}$). For [7, p. 66] $|\tau(\chi)|^{2}=q$, then $\left|\frac{i^{\alpha}q^{\frac{1}{2}}}{\tau(\chi)}\right|=1$. But, at $\epsilon=0$ we have: $1-s=\bar{s}$. Besides, as $\Gamma(\bar{s})=\int_{0}^{\infty}x^{\bar{s}-1}e^{-x}dx=\bar{\Gamma}(s)$ ([8, p. 8]), $\left(\frac{q}{\pi}\right)^{\frac{\bar{s}+\alpha}{2}}=\bar{\left(\frac{q}{\pi}\right)^{\frac{s+\alpha}{2}}}$, and $L(\bar{s},\bar{\chi})=\bar{L}(s,\chi)$ ( (1.1)) , then for 1.2 ), and ( 1.3):
@@ 3. Lemmas on $L(s,\chi_{primitive})$ phase variations along $t$ computed by Euler product
If $\chi$ is not principal, as a primitive character, then $L(s,\chi)$ is an entire function ([3, p. 255] ), and, (1.1) converges absolutely for $\Re(s)>1$, while converges conditionally (see [1, p. 406] ) for $\Re(s)>0$ ( [11, p. 7], and, also unifomily [5, p. 7]). So we take $\Re[s]>0$ throughout.
@@ 3.1. Euler product in critical strip
For example for $p_{max}=31$ we have in (3.2 ) that $N^{o}(terms)=787320$, whoose only 31 belong to $L(s,\chi,p_{max})$. The others belong to $R(s,\chi,p_{max})=R(t,\epsilon,\chi,p_{max})$. While $\rho_{max}(p_{max})$, i.e. the greatest $\rho$ in (3.1), is $Primorial(p_{max})$, so an huge number too, though finite. Besides if we consider a successions of
(3.3) $p_{max}^{k}=\rho_{max}^{k-1}(p_{max}^{k-1})$
it is apparent that all the terms of $R(t,\epsilon,\chi,p_{max}^{k-1})$, are absent in $R(t,\epsilon,\chi,p_{max}^{k})$. So it is apparent that a build-up toward a certain value, from whatever $p^{\prime}_{max}$ to $p^{\prime\prime}_{max}\rightarrow\infty$, it is to be excluded for $R(t,\epsilon,\chi,p^{\prime\prime}_{max})$ when $p^{\prime\prime}_{max}\geq\rho_{max}(p^{\prime}_{max})$. And we are speaking always of huges but finite set of terms.
So we does not need to care, in applying Euler rearrangement for $L(s,\chi,p_{max})=\sum_{n=1}^{p_{max}}\frac{\chi(n)}{n^{s}}$, and $R(s,\chi,p_{max})$, about absolute convergence of the $\prod_{\forall p}\frac{1}{1-\frac{\chi(p)}{p^{s}}}\quad p\quad prime$, because we are dealing with huges, but, finite amount of terms.
@@
It is convenient to think the finite sum in (3.4) as ordered by growing $\rho$ values. We have $p_{max}<\rho<\rho_{max}(p_{max})$. The latter is bigger than $Primorial(p_{max})$, but finite.
@@
Only first term in right part of (3.1), i.e $L(s,\chi,p_{max})=\sum_{n=1}^{p_{max}}\frac{\chi(n)}{n^{s}}$, tend to (C.1). So, considering the whole critical strip, we have:
@@
If we consider the difference (or better the phase difference) in a finite interval, suitably far from $\L (s.\chi)=0$, the conclusion is different. Let us take:
@@
where:
(3.8) $\frac{\partial R(t,\epsilon,\chi,p_{max})}{\partial t}=\sum_{\rho>p_{max}}-\ln(\rho)[\sin(\ln(\rho)t)+i\cos(\ln(\rho)t)]\frac{\chi(\rho)}{\rho^{\Re[s]}}\prod_{\alpha_{j}>0}(1-\chi(p_{j})p_{j}^{-s})^{-1}+$
$\sum_{\rho>p_{max}}[\cos(\ln(\rho)t)-i\sin\cos(\ln(\rho)t)]\frac{\chi(\rho)}{\rho^{\Re[s]}}\times$
$\left\{\sum_{j\ ,\ \alpha(j)\neq 0}-\ln(p_{j})\frac{[\sin(\ln(p_{j})t)+i\cos(\ln(p_{j})t)]}{1-\chi(p_{j})p_{j}^{-s})^{2}}\frac{\chi(p_{j})}{p_{j}^{\Re[s]}}\prod_{\alpha_{k}>0\ k\neq j}(1-\chi(p_{k})p_{k}^{-s})^{-1}\right\}$
In (3.8) we have an huge but finite number of terms with zero mean value ad with phase variation given subtantially by $\ln(\rho)t$. If $p_{max}>>p^{*}$ we have a lot of cancellation in $\Delta t$ integral (3.7) for single $\rho$ (3.8)addend. The cancellation will be complete, for a particular $\rho$, only if
(3.9) $\Delta t\times\ln(\rho)=\frac{2\pi}{\ln(p^{*})}\ln(\rho)=\mu\times 2\pi\ ;\ with\ \mu=m\in N\ ,\ i.e.\ \rho=(p^{*})^{m}$
@@
So the error terms, $\Delta R$, at $\Delta t$ extrema become very small as $p_{max}\rightarrow\infty$, even if extremely slowly, because for different $\rho_{k}$, big enough, we have to sum vectors with different amplitudes but phase almost uniformily distributed in $0,2\pi)$ interval for $t_{1..2}\ln(\rho)$, at least, but here ( differenly then in [15] ), also for the equidistribution of characters , see (3.19 ) .
@@
Let us consider $(R(t,\epsilon,\chi,p_{max})$ terms with $\rho=p_{1}^{\alpha_{1}}...p_{max}^{\alpha_{max}}=p_{max}\times r$, and $\frac{\ln(r)}{\ln(p_{max})}$ small. These terms give an almost null contribution to the differences from $t_{2}$ to $t_{1}$ of (3.1 ), i.e. $L(s,\chi,p_{max})+R(s,\chi,p_{max})$ for $n$ and $\rho$ close to $p_{max}$ because phase difference in $\Delta t$ is: $\ln(\rho)\Delta t=\ln(p_{max})\left(1+\frac{\ln(r)}{\ln(p_{max})}\right)\ \frac{2\pi}{\ln(p_{max})}=2\pi\left(1+\frac{\ln(r)}{\ln(p_{max})}\right)$ , and for $p_{max}\rightarrow\infty$ it tends to $2\pi$. So as concern these, same $\rho$, small $r$ terms we have (considering mean values in $\Delta t$ interval):
@@
where $p^{*}$ is whatever big. The huge amount of terms that are not included in this cancellation (i.e . $\frac{\ln(r)}{\ln(p_{max})}$ not so small , and with $\rho=p_{max}\times r$) have a smaller amplitude and with phases distributed in $0,2\pi)$ interval.
So condition (3.10) is an hint to convergence at lower $p_{max}$ while the argument stemming from identity (3.7) can be applied to generic $p_{max}>>p^{*}$ condition, but, it requires higher $p_{max}$ values to reach a comparable convergence.
In a simplistic view we could say:
(3.4) is mainly an oscillation with period close to $\frac{2\pi}{\ln(p_{max})}$, if we take $\Delta t=m\times\frac{2\pi}{\ln(p_{max})}\ ;\ m\in N$ we filter away the main part of (3.4) for each $\rho=p_{max}\times r$, when $\frac{\ln(r)}{\ln(p_{max})}$ is small. So we have cancellation of main part of $R$.
@@
Conclusions : we must be cautious to affirm that $\L (s.\chi_{primitive})=EP(t,\epsilon,p_{max}\rightarrow\infty)$ in critical strip. Surely where $\L (s.\chi_{primitive})=0$ it is not true as here $R(t,\epsilon,\chi,p_{max})$, i.e. (3.4) is clearly not zero.The sum $R(t,\epsilon,\chi,p_{max})$ contains all the terms of $L(s,\chi)$ from $p_{max}<\rho\leq\rho_{max}$ but the integer containing the primes greater than $p_{max}$ and their powers. If we chose a whatsoever succession of $p^{k}_{max}$, like in (3.3 ) we have that the successive $R(t,\epsilon,\chi,p_{max}^{k})$ must be decorrelated as a noise-like succession, so that a built up is not possible for $p\rightarrow\infty$. It is known that $\sum_{p_{1}}^{\infty}\frac{1}{p^{\gamma}}\rightarrow\infty\ ,\forall p_{1}$, and $\gamma=1$ ( so much more for $\gamma<1$ ). But in (3.4), and, in (3.7), the term $\cos(\ln(\rho)t)-i\sin\cos(\ln(\rho)t)]\frac{\chi(\rho)}{\rho^{\Re[s]}}$ may impose a remarkable cancellation. On the other hand, we are interested in (3.12), i.e. not a simple difference, but, a phase difference. We will see that $|\Delta L|$ is always big meanwhile, for $\Delta t$ fixed, $|\Delta R|$, (3.7), is rather small due to cancellations in bigger $R$ terms by (3.10) choice, and by general cancellation mechanism in interval $\Delta t$, and, we conclude that $\angle[\Delta L+\Delta R]\approx\angle[\Delta L]$ when $\angle[\Delta L]>>0$ ( i.e. close to $s^{*}:L(s^{*},\chi)=0$) or $\frac{\left.\Delta\angle[L(s.\chi)]\right|_{\Delta t}}{\Delta t}>0$ (i.e. on critical line far from zeros, see (A.14 ) , and, Lemma 2). So, where the above conditions are verified, we have some merit to consider :
(3.13) $\frac{\left.\Delta\angle[L(s.\chi)]\right|_{\Delta t}}{\Delta t}=\frac{\left.\Delta\angle[EP(s.\chi)]\right|_{\Delta t}}{\Delta t}\ \ \ \ ;\ \ \Delta t=\frac{2\pi}{\ln(p^{*})}\ ;\ \ p_{max}=(p^{*})^{m};\ m\in N\ ,\ m\rightarrow\infty$
Further examples will confirm this position.
@@ 3.2. Application of Euler Product in critical strip
Expression $\angle[L(s.\chi_{primitive})]=\angle\left[\sum_{n=1}^{\infty}\ \chi(n)/n^{-1}\right]$, where total phase is involved, converges for $\Re(s)>0$, but, in singular points.They are, see (2 ) , $s^{*}_{k}=1/2+\epsilon^{*}_{k}+it_{k}^{*}$,where $\xi(s^{*}_{k},\chi_{primitive})=L(s^{*}_{k},\chi_{primitive})=0$ (see (1.2 ), and (A.2) ). We neglect, for the moment, double zeros. To be specific $\xi(s,\chi)$ is equal ( [7, p. 82]) to the infinite Hadamard product, that has single factors like
@@
Now let us see how to evaluate phase variations of expression $\angle[L(s.\chi_{primitive})]=\angle\left[\sum_{n=1}^{\infty}\ \chi(n)/n^{-1}\right]$ using (3.11 ).
@@
It happens that , for $\{p_{max}\rightarrow\infty\}p^{*}\rightarrow\infty]$, (3.16) grows without bounds not only on the zeros, but also on the segment $0.5<\epsilon<\epsilon^{*}_{k}\ ,\ t=t_{k}^{*}$, where a step discontinuity arises in total phase of (3.14). So, when computing (3.16), we find a kind of ridge stemming from the $k^{th}$ zero $s^{*}_{k}=1/2+\epsilon^{*}_{k}+it_{k}^{*}\ \ ,\ 0.5>\epsilon^{*}_{k}\geq 0$ toward negative reals. Notice that for $p$ big ( i.e. $p^{1/2+\epsilon}>>1\geq\left|\cos(\ln(p)t-\angle[\chi(p)])\right|\ \ \$), the content of braces summed in (3.16 tends to be:
@@
We want to point out at $h$-prime classes in (3.16). If we limit to a finite $p_{max}$ we can write (3.16 ) as the sum of the $\phi(q)$ sums involving the $h$-prime classes:
@@ 3.4. LEMMA 5: Euler Product on critical line
PROOF. From Lemma 2, appendix A and (3.16 ) follows (3.25 ). End of proof.
@@
  • (3) $\frac{\Delta\angle[L_{EP}(t,\epsilon,\chi)]}{\Delta t}$ instead is defined everywhere but points $(\epsilon\leq\epsilon^{*}_{k}\ ,t=t^{*}_{k})$ , see (3.16 ). Here is used total phase.
@@
Suppose $\epsilon>0$ but arbitrarily small, $|t-t^{*}_{k}|>\frac{2\pi}{\ln(p^{*})}$, with (3.6), and $p^{*}$ big (fixed), and $p_{max}\rightarrow\infty$. Then,for (3.25) we have that (3.16) evaluated in $\epsilon=0$ is tending to $\ln\left[\sqrt{\frac{tq}{2\pi}}\right]<0$ as $(p_{max}\rightarrow\infty)p^{*}\rightarrow\infty]$’, in order to verify Lemma 2. We affirm that the same (3.16), evaluated for small $\epsilon>0$ is always negative but with lower absolute value, so that (3.25 ) has a positive increment for small $\epsilon>0$. i,e (at least for odd primitive characters, see (3.24) ):
@@
Negative contribution of (3.16) (and of final result) is given by the $h$-class primes (weighted by (3.29 ) function) inside the analogous intervals $\Delta y_{k,h}$ ( that for $k$ big tends to the previous one $\Delta x_{k,h}$):
@@
Positive contribution of (3.16) (and of final result) is given by the $h$-class primes (weighted by (3.29 ) function) inside the analogous intervals $\Delta y_{k,h}^{\prime}$ ( that for $k$ big tends to the previous one $\Delta x_{k,h}^{\prime}$)
@@
if $m$ is even, the factor $\sin\left(\frac{\pi\ln(x)}{\ln(p^{*})}\right)$ in (3.22) is positive, so, the contributions in corresponding $x$-intervals are as above. In other words, the sign contributions of final result, i.e. (3.16 ) minus (3.22), or , (3.23), in intervals $\Delta x,\Delta x^{\prime},\Delta y,\Delta y^{\prime}$, is as that stated in ( 3.32), (3.35) , and, (3.31), (3.34). Instead where $m$, in (3.36), is odd, then, in every interval above the sign changes concurrently. There are also intermediate situations, but, we will see they can be neglected. So we could compute (3.16) minus (3.22), till $p_{max}$, ( interchanging $O^{+}\leftrightarrow O^{-}$ when $m$ is odd in (3.36) ) as in :
@@
We stress again, (3.37 ) is computed without changing the order given by $p$ or $x=y$, like in (3.21 ) , i.e. partial sums are unaffected. Limit difference in (3.37 ) is in form: $\infty$-$\infty$, and, we know that if $|t-t_{k}^{*}|>\frac{2\pi}{\ln(p^{*})}$ limit result is finite . In other words: positive quantities: $O^{+}_{kh}(\epsilon)_{\int dLi}\ \ ,\ \ O^{-}_{k,h}(\epsilon)_{\int dLi}$,and , $O^{+}_{k,h}(\epsilon)_{\sum}\ \ ,\ \ O^{-}_{k,h}(\epsilon)_{\sum}$ like in ( 3.32), and (3.35) , and, (3.31), (3.34) are only “place-holders” filled according to progress of $x=y<p_{max}$, and with chosen sign following the $m$ parity in (3.36). Consider now the ratios :
@@
for each $h$ and for $k$ big, defined respectively in intervals $\Delta y_{k,h}\rightarrow\Delta x_{k,h}$ ( see (3.31) and (3.32 ) ) , and in intervals $\Delta y_{k,h}^{\prime}\rightarrow\Delta x_{k,h}^{\prime}$ ( see (3.34) and (3.35) ) .
LEMMA 7: $\forall h,k$ as $k\rightarrow\infty$ both (3.40 ) tend to unity.
PROOF. PNT ratio in small intervals [9, p. 23] assures that:
@@
i.e. for $x$ big enough the distribution of primes in equal-sign intevals $\Delta x_{k,h}\rightarrow\Delta y_{k,h}$, and, for (3.19) also $h$-class primes, can be close to theorethical one as we want. In [18, p. 65] it is dealt about the sum of prime sampled functions: “in the neighborhood of the number $x$ the average density of the primes is $1/\ln(x)$. On this basis, if one should wish an estimate for the sum of $F(p)$ over all primes $p<x$, the natural approximation would be”
@@
Now we want the ratio of all positive contributions to final result (3.37 ) at different $\epsilon$ value. In detail : $\forall h<q\ ,gcd(h,q)=1$, $\epsilon^{\prime\prime}>0$, and $\epsilon^{\prime}=0$:
@@
(3.47 ) and (3.46 ), if $\epsilon>0$ is fixed and $p_{max}\rightarrow\infty$ goes to zero, because if $\epsilon>0$ and $k$ i.e. $p_{max}\rightarrow\infty$, then $p^{-\epsilon}\rightarrow 0$. But we have $\forall p_{max}$:
@@
Besides the indeterminate limit (3.37 ), of $\infty$-$\infty$ form, for $\epsilon>0$ ($p^{*}$ is big and fixed), gives :